# Asynchronous Predictive Counterfactual Regret Minimization+ Algorithm in Solving Extensive-Form Games

### v1

@@ 6 Conclusions
$\leq$ $\langle{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I}-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right).$
@@
From the facts that ${\bm{r}}^{t}_{I}=v^{\sigma^{t}}(I)-\langle v^{\sigma^{t}}(I),\sigma^{t}_{i}(I)\rangle\bm{1}$ and $\sigma^{t}_{i}(I)=\frac{[\hat{\bm{\theta}}^{t}_{I}]^{+}}{\|[\hat{\bm{\theta}}^{t}_{I}]^{+}\|_{1}}=\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}$, we have
$\langle-{\bm{r}}^{t}_{I},\hat{\bm{\theta}}^{t}_{I}-\sigma_{i}(I)\rangle=-$ $\langle\langle v^{\sigma^{t}}(I),\sigma^{t}_{i}(I)\rangle\bm{1}-v^{\sigma^{t}}(I),\sigma_{i}(I)-\hat{\bm{\theta}}^{t}_{I}\rangle$ (25)
=-= - $\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-v^{\sigma^{t}}(I),\sigma_{i}(I)-\hat{\bm{\theta}}^{t}_{I}\rangle$
=-= - $\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-v^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-v^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
=-= - $\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-v^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1},\hat{\bm{\theta}}^{t}_{I}\rangle+\langle v^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
=-= - $\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-v^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\|\hat{\bm{\theta}}^{t}_{I}\|_{1}+\langle v^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
=-= - $\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-v^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle v^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle+\langle v^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
=-= - $\langle\langle v^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-v^{\sigma^{t}}(I),\sigma_{i}(I)\rangle$
=-= - $\langle v^{\sigma^{t}}(I),\sigma^{t}_{i}(I)-\sigma_{i}(I)\rangle=\langle v^{\sigma^{t}}(I),\sigma_{i}(I)-\sigma^{t}_{i}(I)\rangle,$
@@
where the fifth line comes from $\hat{\bm{\theta}}^{t}_{I}\geq\bm{0}$, as well as the last line is from $\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}=\sigma^{t}_{i}(I)$ and $\langle\bm{1},\sigma_{i}(I)\rangle=1$ (as $\sigma_{i}(I)\in\Delta^{|A(I)|}$). In addition, we have
$R^{T}(I)=$ $\max_{a\in A(I)}\sum_{t=1}^{T}v^{\sigma^{t}}(Ia)-\sum_{t=1}^{T}\sum_{a\in A(I)}\sigma^{t}_{i}(Ia)v^{\sigma^{t}}(Ia)\leq\max_{\sigma_{i}(I)}\sum_{t=1}^{T}\langle v^{\sigma^{t}}(I),\sigma_{i}(I)-\sigma^{t}_{i}(I)\rangle.$ (26)
Therefore, from Eq. (24), we can bound
@@
$\langle{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right),$ (27)
@@
$\sum^{T}_{t=1}\langle{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\sum^{T}_{t=1}\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+\sum^{T}_{t=1}\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$ (28)
$\leq$ $\frac{1}{\eta}\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})+\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|_{2}\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}+\sum^{T}_{t=1}\frac{\alpha^{t}_{I}}{\eta}\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})+\sum^{T}_{t=1}\frac{1+\alpha^{t}_{I}}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
$\leq$ $\frac{1}{\eta}\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})+\sum^{T}_{t=1}\eta\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{2(1+\alpha^{t}_{I})}+\sum^{T}_{t=1}(1+\alpha^{t}_{I})\frac{\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}}{2\eta}+\sum^{T}_{t=1}\frac{\alpha^{t}_{I}}{\eta}\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})+\sum^{T}_{t=1}\frac{1+\alpha^{t}_{I}}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
@@
where the second inequality comes from that $\forall a,b,\rho>0,ab\leq a^{2}/(2\rho)+\rho b^{2}/2$ ($a=\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|_{2}$, $b=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}$, and $\rho=(1+\alpha^{t}_{I})/\eta$), the third inequality is from $\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|^{2}_{2}/2$ (which is from the fact that in PCFR+ variants, $\psi(\cdot)=\|\cdot\|^{2}_{2}/2$), as well as the last line comes from the facts that $\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})=\|{\bm{\theta}}^{t+1}_{I}-{\bm{\theta}}^{t}_{I}\|^{2}_{2}/2$ and $\eta\bm{R}^{t}_{I}=\bm{\theta}^{t}_{I}$ (see more details in Section D of Farina et al. (2021)). For the term $\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})$ and , we have
@@
where the second equality comes from the definition of ${\bm{\theta}}^{1}_{I}$, i.e., ${\bm{\theta}}^{1}_{I}=\bm{0}$. By substituting Eq. (29) into Eq. (28), we have

### v2

@@ Acknowledgements
$\leq$ $\langle{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I}-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+$
$\quad\quad\quad\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right).$
@@
From the facts that (i) ${\bm{r}}^{t}_{I}=\bm{v}^{\sigma^{t}}(I)-\langle\bm{v}^{\sigma^{t}}(I),\sigma^{t}_{i}(I)\rangle\bm{1}$ and (ii) $\sigma^{t}_{i}(I)=\frac{[\hat{\bm{\theta}}^{t}_{I}]^{+}}{\|[\hat{\bm{\theta}}^{t}_{I}]^{+}\|_{1}}=\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}$, we have
$\langle-{\bm{r}}^{t}_{I},\hat{\bm{\theta}}^{t}_{I}-\sigma_{i}(I)\rangle=$- $\langle\langle\bm{v}^{\sigma^{t}}(I),\sigma^{t}_{i}(I)\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)-\hat{\bm{\theta}}^{t}_{I}\rangle$ (23)
$=$- $\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)-\hat{\bm{\theta}}^{t}_{I}\rangle$
$=$- $\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
$=$- $\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1},\hat{\bm{\theta}}^{t}_{I}\rangle+\langle\bm{v}^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
$=$- $\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\|\hat{\bm{\theta}}^{t}_{I}\|_{1}+\langle\bm{v}^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
$=$- $\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)\rangle-\langle\bm{v}^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle+\langle\bm{v}^{\sigma^{t}}(I),\hat{\bm{\theta}}^{t}_{I}\rangle$
$=$- $\langle\langle\bm{v}^{\sigma^{t}}(I),\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}\rangle\bm{1}-\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)\rangle$
$=$- $\langle\bm{v}^{\sigma^{t}}(I),\sigma^{t}_{i}(I)-\sigma_{i}(I)\rangle=\langle\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)-\sigma^{t}_{i}(I)\rangle,$
@@
where the fifth line comes from $\langle\bm{1},\hat{\bm{\theta}}^{t}_{I}\rangle=\|\hat{\bm{\theta}}^{t}_{I}\|_{1}$, as well as the last line is from $\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}=\sigma^{t}_{i}(I)$ and $\langle\bm{1},\sigma_{i}(I)\rangle=1$ (as $\sigma_{i}(I)\in\Delta^{|A(I)|}$). In addition, we have
$R^{T}(I)=$ $\max_{a\in A(I)}\sum_{t=1}^{T}v^{\sigma^{t}}(Ia)-\sum_{t=1}^{T}\sum_{a\in A(I)}\sigma^{t}_{i}(Ia)v^{\sigma^{t}}(Ia)$ (24)
$\leq$ $\max_{\sigma_{i}(I)}\sum_{t=1}^{T}\langle\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)-\sigma^{t}_{i}(I)\rangle.$
Therefore, from Eq. (22) and $\langle-{\bm{r}}^{t}_{I},\hat{\bm{\theta}}^{t}_{I}-\sigma_{i}(I)\rangle=\langle\bm{v}^{\sigma^{t}}(I),\sigma_{i}(I)-\sigma^{t}_{i}(I)\rangle$, we can bound
@@
$\langle{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+$ (25)
$\quad\quad\quad\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right),$
@@
$\sum^{T}_{t=1}\langle{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\rangle+\sum^{T}_{t=1}\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+$ (26)
$\quad\quad\quad\sum^{T}_{t=1}\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
$\leq$ $\frac{1}{\eta}\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})+\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|_{2}\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}+$
$\quad\quad\quad\sum^{T}_{t=1}\frac{\alpha^{t}_{I}}{\eta}\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})+\sum^{T}_{t=1}\frac{1+\alpha^{t}_{I}}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
$\leq$ $\frac{1}{\eta}\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})+\sum^{T}_{t=1}\eta\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{2(1+\alpha^{t}_{I})}+\sum^{T}_{t=1}(1+\alpha^{t}_{I})\frac{\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}}{2\eta}+$
$\quad\quad\quad\sum^{T}_{t=1}\frac{\alpha^{t}_{I}}{\eta}\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})+\sum^{T}_{t=1}\frac{1+\alpha^{t}_{I}}{\eta}\left(-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$
@@
where the second inequality comes from that $\forall b,c,\rho>0,bc\leq b^{2}/(2\rho)+\rho c^{2}/2$ ($b=\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|_{2}$, $c=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|_{2}$, and $\rho=(1+\alpha^{t}_{I})/\eta$), the third inequality is from $\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})=\|{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}\|^{2}_{2}/2$ (which is from the fact that in PCFR+ variants, $\psi(\cdot)=\|\cdot\|^{2}_{2}/2$), as well as the last line comes from the facts that $\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})=\|{\bm{\theta}}^{t+1}_{I}-{\bm{\theta}}^{t}_{I}\|^{2}_{2}/2$ and $\eta\bm{R}^{t}_{I}=\bm{\theta}^{t}_{I}$ (see more details in Section D of farina2021faster). For the term $\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{1}_{I})$ and , we have
@@
where the second equality comes from the definition of ${\bm{\theta}}^{1}_{I}$, i.e., ${\bm{\theta}}^{1}_{I}=\bm{0}$, and the inequality is from $\|\sigma_{i}(I)\|_{1}\leq 1,\forall\sigma_{i}(I)\in\Delta^{|A(I)|}$. By substituting Eq. (27) into Eq. (26), we have